OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP WILL FAN AND MEDAL!!! A country's population in 1991 was 136 million. In 2000 it was 141 million. Estimate the population in 2016 using the exponintial growth formula. Round your answer to the nearest million. P=Ae^kt

OpenStudy (mathmale):

I'd suggest that you let 1 represent the year 1991, 2 1992, and so on, up through 16 to represent 2016. Which integer would then represent the year 2000?

OpenStudy (mathmale):

In \[A=Pe ^{kt}\]

OpenStudy (anonymous):

No. I need the actual number lol one sec. I asked a question like this yesterday but I still am not good at doing them,

OpenStudy (mathmale):

you have two constants: One is P, the original population. The other is k, which is called the "growth constant."

OpenStudy (mathmale):

Can you use the given info to determine P and ki?

OpenStudy (er.mohd.amir):

here P ,A,k,t are value know to solve tis problem

OpenStudy (anonymous):

One second OS is lagging ><

OpenStudy (er.mohd.amir):

in problem U wants to find P if t is given but how u find A and k for A and k use given data

OpenStudy (anonymous):

You just confused me....

OpenStudy (mathmale):

Ask some questions, then. You seem to have experienced a similar math problem in the near past; what could you apply from that experience towards solving this new problem?

OpenStudy (mathmale):

Again, the function in question is\[A=Pe ^{kt}\]

OpenStudy (anonymous):

But if you just said P is the original population and @Er.Mohd.AMIR aid we're solving for p.

OpenStudy (mathmale):

Yes, I see how my comments could have confused you. The formula for population is\[P=Ae ^{kt}\]

OpenStudy (mathmale):

You are told that in the year 1991 (which I recommend denoting by t=1), the initial population, A, was 136 (representing 136 million).

OpenStudy (mathmale):

You are also told that in 2000 (represented by 9), the population, P, is 141 (for 141 million). You have two sets of facts, the info necessary for you to find the values A and k. Once you've done that, you need only substitute t=16 into your equation for the population, P, for the year 2016 (represented by t=16).

OpenStudy (anonymous):

where are you getting 9?

OpenStudy (anonymous):

There's no9 or representing in this...

OpenStudy (mathmale):

1 = t represents the year 1991, 2 reps. 1992, .... 8 reps. 1998 9 reps 1999 10 reps 2000. Sorry. My bad.

OpenStudy (anonymous):

But you don't do that for this? I've done it before and though I don't know how to fully do it I know you don't do that.....

OpenStudy (mathmale):

I'm sorry you don't care for my approach. Choosing 1 to represent 1991, etc., is a common approach to solving this type of problem. There would be NO advantage to your using 1991, 1992, etc., instead of my suggested 1, 2, 3, etc. If you want to approach this problem solution in another way, perhaps you'd be better off waiting for another user to help you.

OpenStudy (anonymous):

:/ I just don't get that.. it seems like you're pulling numbers out of the air.

OpenStudy (er.mohd.amir):

if u use scientefic calculator u can use given values

OpenStudy (anonymous):

I don't have one @Er.Mohd.AMIR

OpenStudy (mathmale):

You'll either need to obtain a suitable calculator (I recommend the latest version of the TI-84), or you'll need to use online calculators. This problem is not one to do with pencil and paper.

OpenStudy (er.mohd.amir):

simply put t at 1991 is t=1 and p=136 mil. and for 2000 t=10 since diff in years is 10 and p=141 u get two equations in A and k

OpenStudy (mathmale):

I still suggest you should let 1 (not 0) represent 1991. Right: doing this properly should result in two separate equations for A and k that can be solved simultaneously for those two constants.

OpenStudy (er.mohd.amir):

where are u?

OpenStudy (mathmale):

@whateven: sorry, but I need to get off the 'Net for now. Good luck. Try to get your hands on an appropriate calculator, and review what you did yesterday in a similar problem solving situation.

OpenStudy (anonymous):

I had to do kaundry and there's only one p...

OpenStudy (er.mohd.amir):

136=A e^(1*k) and 141=A e^(10*k)

OpenStudy (er.mohd.amir):

first divide these eq.

OpenStudy (anonymous):

Ohhhh ok.

OpenStudy (anonymous):

but what is e

OpenStudy (er.mohd.amir):

e=2.718 its Eular no.or exponential no.

OpenStudy (anonymous):

will it always be 2.718

OpenStudy (er.mohd.amir):

yes

OpenStudy (er.mohd.amir):

e=2.718281828..............................................

OpenStudy (er.mohd.amir):

its non teminated no.

OpenStudy (anonymous):

So i divide e by 1 and 10?

OpenStudy (er.mohd.amir):

141/136=e^(9k)

OpenStudy (er.mohd.amir):

taking log both sides wit base e to solve it

OpenStudy (mathmale):

"e" is the most common base used in exponential functions. The calculator you use for these calculations will ideally have a key for "e." You do not need to memorize the value of e for now, but if you're curious, yes, the value is around 2.718. Once again: You should end up with two separate equations in A and k, and solve these two equations simultaneously for A and k. Then write out \[P=Ae ^{kt}\]. Substitute t=16 (or whatever) for t, to calculate the probable population in the year 2016.

OpenStudy (anonymous):

(am eating almost done sorry. where is this 9 coming from?)

OpenStudy (er.mohd.amir):

e^10k / e^k= e^(10k-k)= e^9k

OpenStudy (mathmale):

You've asked that question before. I was wrong there; t=10 better represents the year 2000. See our previous discussion, above.

OpenStudy (er.mohd.amir):

if base is same then power is subtract in division

OpenStudy (anonymous):

oh ok. so what do we do after that?

OpenStudy (er.mohd.amir):

u have log in calculator use that

OpenStudy (anonymous):

Do you know of an online calculator that will do that?

OpenStudy (er.mohd.amir):

which OS u use in window a calculator avialable

OpenStudy (anonymous):

What?

OpenStudy (er.mohd.amir):

log (141/136)=9k log e 9 k =0.036105004 k=0.004011667182 since log e=1

OpenStudy (er.mohd.amir):

use k=0.004

OpenStudy (er.mohd.amir):

now 136=A* e^k put k=0.004 get A

OpenStudy (er.mohd.amir):

@WhatEven finish it quickly

OpenStudy (anonymous):

Am here os is lagging

OpenStudy (er.mohd.amir):

e^0.004=1.004 so A=136/1.004=135.4570

OpenStudy (anonymous):

so then 135? :)

OpenStudy (er.mohd.amir):

its A=135.45 and k=0.004 u get k and A noe P=A e^kt put t=16 get ans

OpenStudy (er.mohd.amir):

sorry t=26

OpenStudy (anonymous):

so then \[p=135.45*e ^{k26}\]

OpenStudy (er.mohd.amir):

put value of k also

OpenStudy (er.mohd.amir):

e^(0.004*26)=e^(0.104)=1.1096

OpenStudy (anonymous):

\[p=134.45*e ^{.004*26}\]

OpenStudy (er.mohd.amir):

P=135.45*1.1096=150.295

OpenStudy (er.mohd.amir):

fine

OpenStudy (anonymous):

fine?

OpenStudy (er.mohd.amir):

? what

OpenStudy (anonymous):

Why did you say fine o.o

OpenStudy (er.mohd.amir):

if u have good calculator thhen problem is easy

OpenStudy (anonymous):

I don't have one. My calculator is broken

OpenStudy (er.mohd.amir):

fine is for me when i complete something.

OpenStudy (er.mohd.amir):

ok i m now turnoff from here.

OpenStudy (anonymous):

...I don't ave a way to calculate that last part though...

OpenStudy (er.mohd.amir):

take cal. of any one and solve yourself

OpenStudy (anonymous):

what?

OpenStudy (er.mohd.amir):

use calculaor of any of Ur friend for solving it urself

OpenStudy (anonymous):

I do not have a calculator.

OpenStudy (anonymous):

I am also home schooled no friends to ask that's the whole reason I use open study if I were in a brick and mortar school I could just ask the teacher

OpenStudy (er.mohd.amir):

use google and write online calaulator

OpenStudy (anonymous):

(Waiting on it to load ._.

OpenStudy (anonymous):

wait so 150?